OpenStudy (anonymous):

how do i get z from 11z-z=70?

OpenStudy (anonymous):

11z -1z = 70 10z = 70 so z = 7

OpenStudy (anonymous):

11z-z=70 11z-1z=70 10z=70 x=7 you need to try these on your own there very easy

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

now i grt it thank you relly much

OpenStudy (anonymous):

np